Hi there,
I am not sure how I should solve this problem,can someone maybe explain it to me?
3.5x-1=-1/2(-7x+2)
Thank you

Answer by Alan3354(69443)

3.5x-1=-1/2(-7x+2)
---
If you mean
3.5x-1=-(1/2)*(-7x+2)
Multiply by 2 to eliminate fractions
7x - 2 = -(-7x+2) = 7x - 2
----
No unique solution, equal for all values of x.